OpenStudy (anonymous):

Select inductive reasoning, deductive reasoning, or neither. If 5x + 7 = 12, then x = 1. Inductive reasoning Deductive reasoning neither

OpenStudy (zzr0ck3r):

deductive: assume something then get to the conclusion from logical steps. inductive: get to the hypothesis from specific examples neither: neither

OpenStudy (zzr0ck3r):

which one do you think it is?

OpenStudy (zzr0ck3r):

examples: deductive: all men were boys, john is a man and thus at one point was a boy inductive: cows give milk, goats give milk, thus all 4 legged animals give milk neither: random guess

OpenStudy (zzr0ck3r):

what you have assume 5x+7=12 then from what we know of algebra 5x+7-7=12-7 so 5x=5 so x=1 we assumed 5x+7=12 and got to the conclusion that x must be 1

OpenStudy (zzr0ck3r):

@miadeex ? what sort of reasoning is that^^^

OpenStudy (zzr0ck3r):

@miadeex ? please answer since im taking the time to help.

OpenStudy (anonymous):

uhmm deductive ?

OpenStudy (zzr0ck3r):

correct